A study conducted at a certain college shows that 62% of the schools graduates find a job in their chosen feild within a year after graduation what is the probability that among 5 randomly selected graduates, at least one finds a job in his or her chosen feild
Answer by stanbon(75887)

A study conducted at a certain college shows that 62% of the schools graduates find a job in their chosen feild within a year after graduation what is the probability that among 5 randomly selected graduates, at least one finds a job in his or her chosen field.
----
P(at least one finds a job) = 1 - P(none find a job)
= 1 - 0.38^5
= 1 - 0.007924
= 0.9921
============
Cheers,
Stan H.